OpenStudy (aytek):

find the general formula of (dy/dx)+((2x+1)/x)*y=e^(-2x). y(1)=1/(e^2) (i used ' ^ ' as power sign.)

zepdrix (zepdrix):

\[\large\rm y'+\left(\frac{2x+1}{x}\right)~y=e^{-2x},\qquad\qquad y(1)=\frac{1}{e^2}\]

zepdrix (zepdrix):

First order, linear, so you'll want to find an integrating factor, ya? :)

OpenStudy (aytek):

i want to find y(x) but yes we will use integrating factor

OpenStudy (anonymous):

I.F\[=e ^{ \int\limits \frac{ 2x+1 }{ x }dx}=?\]

OpenStudy (aytek):

i want to find the general solution of y(x)

zepdrix (zepdrix):

I would probably recommend simplifying this fraction before integrating :)\[\large\rm \frac{2x+1}{x}\qquad=\frac{2x}{x}+\frac{1}{x}\qquad=2+\frac{1}{x}\]That's a little easier to integrate, ya?

OpenStudy (aytek):

i already did that

OpenStudy (aytek):

i also found a general solution but it does not satisfy y(1)=1/(e^2)

zepdrix (zepdrix):

So what do you get for your integrating factor? :d

OpenStudy (aytek):

e^(2x) * x

zepdrix (zepdrix):

Ah good good good. Multiplying through by our integrating factor,\[\large\rm x e^{2x}y'+e^{2x}(2x+1)y=1\]I simplied already, hopefully you understood that, otherwise I can back up a step. (Don't forget to give the RIGHT SIDE your integrating factor as well)

zepdrix (zepdrix):

Assuming our integrating factor is correct, left side of the equation should reverse product rule, and sure enough, it does here :)\[\large\rm \left(yxe^{2x}\right)'=x\]Woops I messed up the right side on the previous post, sorry about that.

OpenStudy (anonymous):

\[C.S.~is ~y*xe^{2x}=\int\limits e ^{-2x}x e ^{2x}dx=\frac{ x^2 }{ 2 }+c\] when x=1 \[y=\frac{ 1 }{ e^2 }\] \[\frac{ 1 }{ e^2 }*1*e^2=\frac{ 1^2 }{ 2 }+c,c=1-\frac{ 1 }{ 2 }=\frac{ 1 }{ 2 }\] substitute the value of c and find the general solution.

OpenStudy (aytek):

ok, thank you very much

OpenStudy (anonymous):

yw